a man is now four times as old as his son and in four years from now, the product of their ages will be 520. Find their ages now.
Answer by checkley77(12569)

LET THE MAN'S AGE=4X
LET THE SON'S AGE-X
(4X+4)(X+4)=520
4X^2+4X+16X+16=520
4X^2+20X+16-520=0
4X^2+20X-504=0
4(X^2+5X-126)=0
4(X-9)(X+14)=0
X-9=0
X=9 AGE OF THE SON.
4*9=36 AGE OF THE DAD.
PROOF:
(4*9+4)(9+4)=520
(36+4)(13)=520
40*13=520